How many seven-digit telephone numbers are possible if the first two digits cannot be ones or zeros?
Answer by Edwin McCravy(20081)

Choose the 1st digit any of 8 ways, {2,3,4,5,6,7,8,9}.
Choose the 2nd digit any of 8 ways, {2,3,4,5,6,7,8,9}.
Choose the 3rd digit any of 10 ways, {0,1,2,3,4,5,6,7,8,9}.
Choose the 4th digit any of 10 ways, {0,1,2,3,4,5,6,7,8,9}.
Choose the 5th digit any of 10 ways, {0,1,2,3,4,5,6,7,8,9}.
Choose the 6th digit any of 10 ways, {0,1,2,3,4,5,6,7,8,9}.
Choose the 7th digit any of 10 ways, {0,1,2,3,4,5,6,7,8,9}.

Answer 8*8*10*10*10*10*10 = 82*105 = 6400000

Edwin